A support group has 14 patients that attend each week. Suppose 3 patients are selected at random to participate in a study. How many different samples of this size can be selected from this population of 14 patients using experimental sampling and theoretical sampling?

(a) experimental sampling
samples

(b) theoretical sampling
samples
Answer #1

Answer:

Given that:

A support group has 14 patients that attend each week. Suppose 3 patients are selected at random to participate in a study. How many different samples of this size can be selected from this population of 14 patients using experimental sampling and theoretical sampling?

a) Experimental sampling:

Numbr of samples possibe=NCn

N=14,n=3

Number of samples possible=14C3=14!/3!12!

=(14*13*12)/(3*2*1)

= 364

b)Theoretical sampling:

Number of possible sample=Nn= 14^3 = 2744